Write the equation of the line perpendicular to the line given by 3x – 2y=9 and having the same y intercepts as the line given by 2x + 5y = 0

Answer by josgarithmetic(39617)

3x-2y=0 will have the same y (and x) intercept as 2x+5y=0; but since you want the line PERPENDICULAR to 3x-2y=9, the equation you want is 2x+3y=0.

Why so quick showing no steps?
Ax+By=C shows slope information in the coefficients A and B. Try solving this standard form equation for y and adjust to slope-intercept form and this will be more clear.